OpenStudy (anonymous):

Use the commutative law of addition to write an equivalent expression. 6x + 5y

OpenStudy (turingtest):

commutative law of addition is\[a+b=b+a\]

OpenStudy (anonymous):

ok still lost

OpenStudy (turingtest):

call one term in your problem 'a' and the other 'b' then apply the rule

OpenStudy (anonymous):

so 6+x = 5+y?

OpenStudy (turingtest):

not quite let a=6x let b=5y not apply the rule

OpenStudy (turingtest):

now apply...*

OpenStudy (anonymous):

so 6x+5y=5y+6x

OpenStudy (turingtest):

yes